There are two asset classes: stocks and bonds. The expected return to stocks is 12% with...

  1. There are two asset classes: stocks and bonds. The expected return to stocks is 12% with a standard deviation of 22%. The expected return to bonds is 5% with a standard deviation of 8%. The correlation between stocks and bonds is 0.80. Assume your utility function is given by: U = E(r) – 2.5σ2. Given no constraints on investing in stocks or bonds, your utility is highest in which of the following portfolios?
Answer #1

Solution

Given : return from stock =12%

Standard deviation of stock = 22%

Return from bonds = 5%

Standard deviation of bond = 8%

utility = E(r) -2.5 variance

1) when only stock in the portfolio

Utility = E(r) - 2.5 variance

= 0.12 - 2.5 ( 0.22)^2

= 0.12 - 2.5 (0.0484)

= 0.12 - 0.121

= -0.001 or -0.1%

2) when only bond is used in portfolio

Utility = E(r) - 2.5 variance

= 0.05 - 2.5 (0.08)^2

= 0.05 - 2.5 (0.0064)

= 0.05 - 0.016

= 0.034 or 3.4%

Therefore the utility is highest when the portfolio consists only of bonds.